MTA Monitoring Report 09/09/05 - 09/15/05

Archive of all weekly reports

Archive of monthly reports

An ACIS DEA power off anomaly occured at about 258:23:30. Investigation underway. Load SEP1505C is running and executing target maneuvers, but ACIS is not taking data.


Several radiation alerts (ACE, GOES, Kp) were issued on Sep 8 and Sep 15.
SCS-107 ran at day 257:0106z due to high radiation.

Lost science time:
9/8-9/13 shutdown: 285 ks
9/14-9/15 shutdown: 80 ks

Detrended CTI

CTI page was updated. The top panel shows an example of detrended CTI measurements from the front illuminated CCD3 Node 0. The top plot is all data points without any temperature dependency correction. The middle plot is data with the focal temperature equal or colder than -119.7 C, and the integration time longer than 7,000 sec. The bottom plot is all data with the temperature correction.

The values quoted below are temperature factor corrected ones.
Average of ACIS-I CCDs (MnKa) slope 4.378e-9 CTI/day
Detrended slope 7.696e-9 CTI/day

ACIS Warm Pixels

We display all bad pixels showed up in past, and, if any, previously unknown bad pixels appeared in the last 14 days. We also list hot pixels (defined as 1000 above the average of bias background). Warm columns 509-514 are probably due to a computational artificial effect due to boundaries.

ACIS Focal Plane Temperature

SIM Movements

4 TSC moves this period
weekly average time/step 0.00169 s
mission average time/step 0.00136 s



Recent Observations

"Prob" means that there was probably bad time data included in the manual runs. This will probably be corrected in future AP processing.

5555 ACIS-S HETG The Crab Nebula OK /Grat OK


This Week's focus is ACIS.

Trends last reported on June 23.

Only the most interesting or representative msids are shown below. For a full listing choose the bulletted link.